The sum of all odd integers between 2 and 100 divisible by 3 is
(a) 17
(b) 867
(c) 876
(d) 786

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Answer: b Explaination:Reason: The numbers are 3, 9,15, 21, …, 99 Here a = 3, d = 6 and an = 99 ∴ an = a + (n – 1 )d ⇒ 99 = 3 + (n – 1) x 6 ⇒ 99 = 3 + 6n – 6 ⇒ 6n = 102 ⇒ n = 17